You have the following demand for product X: Q = 200 - 0.30P. At a price =$100, point price elasticity of demand is about

A) -1.50

B) -0.176

C) can't be determined since the quantity demanded at p = $100 is not known

D) -1.176
Answer #1

B) -0.176

Explanation: Following is the formula for point elasticity:

dQ/dP * P/Q

dQ/dP = d/dP (200 - 0.30P) = -0.30

When price is 100, Q = 200 - 0.30P = 200 - 0.30*100 = 200 - 30 = 170.

So, P/Q = 100/170 = 0.588

point elasticity = dQ/dP * P/Q = -0.30*0.588 = -0.176